# Activation thresholds and expressiveness of polynomial neural networks

### v1

@@ 3 Equi-width setting
The advantage of depth for neural networks is well-established through both empirical observations and theoretical results [24, 31, 23, 29]. Width also plays a significant role in network expressivity and improves the behavior of the loss surface by smoothing the optimization landscape of the loss function [18, 27, 28]. The study of networks with bounded width appears in the context of universal approximation theorems [17, 19, 22] as well.
For sufficiently high activation degree, a neurovariety has the expected dimension by Theorem 12. For specific architectures, better bounds on the activation threshold may be computed. In this section, we focus on determining the expressivity of polynomial neural networks with equi-width architecture, meaning the layers have equal widths, i.e., $d_{0}=d_{1}=\dots=d_{L}$ in Definition 1.
We begin with a proposition which will be utilized in the main result of this section.
@@
Let $\mathbf{d}=(d_{0},\dotsc,d_{L})$ be an equi-width architecture such that $d:=d_{0}=d_{1}=\dotsb=d_{L}$ and let $\sigma_{i}:\mathbb{R}^{d_{i}}\to\mathbb{R}^{d_{i}}$ be the activation map given by the coordinate-wise exponentiation to the $r$-th power. A polynomial neural network
@@
If some $W_{i}$ is singular, then let $i_{0}$ be the minimal index such that $W_{i_{0}}$ is singular and $v\in\ker W_{i_{0}}$ be a non-zero vector. For each $j<i_{0}$, the linear map $W_{j}$ is surjective, and each $\sigma_{j}$ is surjective as well. Thus, there exists non-zero $x_{\ast}\in\mathbb{C}^{d}$ such that
@@
We now provide the main result of this section, that the activation threshold of an equi-width architecture is one. That is, for an equi-width architecture $\mathbf{d}$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r\geq 2$.
This is done by explicitly computing a single fiber of the parameter map $\Psi_{\mathbf{d},r}$. This dimension provides an upper bound on the generic dimension of a fiber, and so we obtain a lower bound on the dimension $\dim\mathcal{V}_{\mathbf{d},r}$. This lower bound is in fact equal to our upper bound given by the expected dimension, thereby proving the result.
@@
If $\mathbf{d}=(d_{0},\dotsc,d_{L})$ is an equi-width architecture with $d=d_{0}=\dotsb=d_{L}$, $L>1$ and $d>1$, then the activation threshold is $\operatorname{ActThr}(\mathbf{d})=1$. That is, if $\mathbf{d}$ is equi-width with $L>1$ and $d>1$, then for all $r>1$ the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension
@@
Proof.
[…]
We remark that when $L>1$, $\operatorname{ActThr}(\mathbf{d})>0$. Indeed, if $r=1$, then $\mathcal{V}_{\mathbf{d},r}$ is again the space of linear functions from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$ so that $\mathcal{V}_{\mathbf{d},r}=d^{2}$. Then $d>1$ implies that
$\dim\mathcal{V}_{\mathbf{d},r}=d^{2}<Ld^{2}-(L-1)d=\text{edim}\mathcal{V}_{\mathbf{d},r}.$
We show that when $L>1$ and $r>1$, a specific fiber of the parameter map
$\Psi_{\mathcal{V}_{\mathbf{d},r}}:\mathbb{R}^{d_{L}\times d_{L-1}}\times\dotsb\times\mathbb{R}^{d_{1}\times d_{0}}\to\mathcal{V}_{\mathbf{d},r}$
has dimension $(L-1)d$ and thus $\operatorname{ActThr}(\mathbf{d})=1$. Consider the neural network
$\begin{pmatrix}x_{1}^{r^{L-1}}\\ \vdots\\ x_{d}^{r^{L-1}}\end{pmatrix}=\Psi_{\mathbf{d},r}(I_{d},\dotsc,I_{d}).$
Precisely, we show that the fiber $\Psi_{\mathbf{d},r}^{-1}(x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}})$ is the set of multi-homogeneities given in Lemma 5.
To this end, let $W_{1},\dotsc,W_{L}$ be weight matrices such that
We will show by induction that the output of the $i$-th layer consists of scalar multiples of the monomials $x_{1}^{r^{i-1}},\dotsc,x_{d}^{r^{i-1}}$, which will imply that each $W_{i}$ is a (invertible) scaled permutation matrix. Consider the output of the $(L-1)$-th layer,
$\begin{pmatrix}p_{\mathbf{w}_{1}}(x)\\ \vdots\\ p_{\mathbf{w}_{d}}(x)\end{pmatrix}=(W_{L-1}\sigma_{L-2}W_{L-2}\dotsb W_{2}\sigma_{1}W_{1})(x).$
There is then an equality
$W_{L}\begin{pmatrix}p_{\mathbf{w}_{1}}^{r}(x)\\ \vdots\\ p_{\mathbf{w}_{d}}^{r}(x)\end{pmatrix}=\begin{pmatrix}x_{1}^{r^{L-1}}\\ \vdots\\ x_{d}^{r^{L-1}}\end{pmatrix}$ (7)
Since the polynomial neural network $(x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}})$ only has the trivial zero, Proposition 13 implies that $W_{L}$ is invertible. Multiplying both sides of (7) by $W_{L}^{-1}$, each $p_{\mathbf{w}_{i}}^{r}(x)$ is a linear combination of $x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}}$. For $r>1$, this is only possible if each $p_{\mathbf{w}_{1}}$ is a monomial. Indeed, no linear combination of $x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}}$ is a pure $r$-th power except for scalar multiples of these individual monomials. Thus each $p_{\mathbf{w}_{i}}^{r}$ is a scalar multiple of some $x_{j}^{r^{L-1}}$ and each $p_{\mathbf{w}_{i}}$ is a scalar multiple of some $x_{j}^{r^{L-2}}$.
It follows that $W_{L}$ is a scaled permutation matrix—it is an invertible matrix that sends a set of monomials to a set of scalar multiples of those monomials. As the output of the $(L-1)$-th layer is then a set of scalar multiples of $x_{i}^{r^{L-2}}$, we inductively apply this argument to show that the fiber has the form of Lemma 5.
The fiber $\Psi_{\mathbf{d},r}^{-1}(\Psi_{\mathbf{d},r}(I_{d},\dotsc,I_{d}))$ then consists of the set of multi-homogeneities and has the expected dimension $\sum_{j=1}^{L-1}d_{j}$, which is an upper bound on the generic dimension of a fiber. It follows that
$\text{edim}\mathcal{V}_{\mathbf{d},r}=\sum_{i=0}^{L-1}d_{i}d_{i+1}-\sum_{j=1}^{L-1}d_{j}$
is a lower bound for $\dim\mathcal{V}_{\mathbf{d},r}$. As $\text{edim}\mathcal{V}_{\mathbf{d},r}$ is also an upper bound, $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r>1$. ∎
@@ 4 Outlook
Bounding the activation threshold
Our bound on the activation threshold for polynomial neural networks is not universally optimal. We expect that a lemma of [15] used to prove Picard theorems for holomorphic maps into Fermat hypersurfaces will allow us to improve the bound in Theorem 12. Effective methods in algebraic geometry to compute the activation threshold of a given architecture also deserve to be studied. One possible approach is to formalize a connection to patterns of linear dependence among families of polynomials. Such patterns have notable examples from number theory independently due to A. H. Desboves [12] and Noam Elkies [10] and were studied in [30] for a finite set of arbitrary polynomials in terms of tickets. The ticket $\mathcal{T}(f)$ of a finite set of polynomials $F=\{f_{j}\}$ is defined as
$T(F)=\{m\in\mathbb{N}:\{f_{j}^{m}\}\text{ is linearly dependent}\}.$
By viewing the polynomial neural network recursively, layer by layer, it should be possible prove results about the expected dimension by translating tickets into the polynomial neural network setting.
Non-increasing architectures
One future direction in this research area is to determine the activation thresholds for non-increasing architectures $\mathbf{d}=(d_{0},\dotsc,d_{L})$ — those architectures satisfying $d_{0}\geq d_{1}\geq\dotsb\geq d_{L}$. By generalizing the proof of Theorem 14, one might hope to prove a conjecture of [21] that for a non-increasing architecture $\mathbf{d}$ with $d_{L}>1$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r>1$. That is, prove that a non-increasing architecture $\mathbf{d}$ has activation threshold less than or equal to one.
@@
Another research direction is to introduce the notion of activation thresholds for other linear and non-linear neural networks. Permitting negative exponents in a polynomial activation map leads to the notion of rational neural networks [5]. There is experimental evidence that it is beneficial to choose rational activation maps that are very low degree. Defining activation thresholds for rational networks would provide a quantitative measure to make these observations rigorous. Rational neural networks are currently being investigated by Alexandros Grosdos et al. Certain linear networks may also admit the definition of an activation threshold as a means to study expressiveness. Tropical rational neural networks provide a lens to study networks with piecewise linear activation maps [3, 8, 9, 33]. Such networks encompass ReLU networks; a ReLU network of fixed architecture is described by a semialgebraic set inside the parameter space of tropical rational functions. In this context, tropical rational functions provide a means to define activation thresholds to frame and obtain results on expressiveness. Moreover, because the dimension of the neurovariety is preserved under tropicalization, activation thresholds may aid in translating expressiveness results for tropical neural networks to polynomial networks.

### v3

@@ 3 Equi-width setting
The advantage of depth for neural networks is well-established through both empirical observations and theoretical results [32, 41, 31, 37]. Width also plays a significant role in network expressiveness and improves the behavior of the loss surface by smoothing the optimization landscape of the loss function [20, 35, 36]. The study of networks with bounded width appears in the context of universal approximation theorems [19, 24, 29] as well.
For sufficiently high activation degree, a neurovariety has the expected dimension by Theorem 18. For specific architectures, better bounds on the activation degree threshold may be computed. In this section, we focus on determining the expressiveness of polynomial neural networks with equi-width architecture, meaning the layers have equal widths, that is, $d_{0}=d_{1}=\dots=d_{L}$ in Definition 1.
We begin with a proposition which will be used in the main result of this section.
@@
Let $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ be an equi-width architecture such that $d:=d_{0}=d_{1}=\dotsb=d_{L}$ and let $\sigma_{i}:\mathbb{C}^{d_{i}}\to\mathbb{C}^{d_{i}}$ be the activation map given by coordinate-wise exponentiation to the $r$-th power. A polynomial neural network
@@
If some $W_{i}$ is singular, then let $i_{0}$ be the minimal index such that $W_{i_{0}}$ is singular and let $v\in\ker W_{i_{0}}$ be a non-zero vector. For each $j<i_{0}$, the linear map $W_{j}$ is surjective, and each $\sigma_{j}$ is surjective as well. Thus, there exists non-zero $x_{\ast}\in\mathbb{C}^{d}$ such that
@@
We now provide the main result of this section, that the activation degree threshold of an equi-width architecture is one. This is done by explicitly computing a single fiber of the parameter map $\Psi_{\mathbf{d},r}$. The dimension of this fiber provides an upper bound on the generic dimension of a fiber, and so we obtain a lower bound on the neurovariety dimension. This lower bound is in fact equal to the upper bound given by the expected dimension, thereby proving the result.
@@
If $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ is an equi-width architecture with $d=d_{0}=d_{1}\dotsb=d_{L}$, $L>1$, and $d>1$, then the activation degree threshold is $\operatorname{ActThr}(\mathbf{d})=1$. That is, if $\mathbf{d}$ is equi-width with $L>1$ and $d>1$, then for all $r>1$ the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension
@@
Proof.
If $r=1$, then $\mathcal{V}_{\mathbf{d},1}$ is the space of linear functions from $\mathbb{C}^{d}$ to $\mathbb{C}^{d}$. Therefore $\dim\mathcal{V}_{\mathbf{d},1}=d^{2}$. Since
$\operatorname{edim}\mathcal{V}_{\mathbf{d},1}=Ld^{2}-(L-1)d,$
we have $\operatorname{defect}(\mathcal{V}_{\mathbf{d},1})=(L-1)(d^{2}-d)$. Therefore for all $L>1$, $\operatorname{ActThr}(\mathbf{d})>0$.
Recall the parameter map (3) for polynomial neural networks. For notational convenience, we let $S_{L,d,r}$ denote the set of parameters that are mapped to $\Psi_{\mathbf{d},r}(I_{d},\dots,I_{d})$:
$S_{L,d,r}:=\left\{(W_{1},\dots,W_{L})\in\left(\mathbb{C}^{d\times d}\right)^{L}:\Psi_{\mathbf{d},r}(I_{d},\dots,I_{d})=\Psi_{\mathbf{d},r}(W_{1},\dots,W_{L})\right\}.$ (16)
We want to show for $L>1$ and $d>1$ that
$\begin{split}S_{L,d,r}=\Big{\{}(W_{1},\dots,\,PDW_{L-1},D^{-r}P^{T})\in\left(\mathbb{C}^{d\times d}\right)^{L}:(W_{1},\dots,W_{L-1})\in S_{L-1,d,r},&\\ \,P\text{ is a permutation matrix and }D\text{ is an invertible diagonal matrix}\Big{\}}.\end{split}$ (17)
Note that if $(W_{1},W_{2},\dotsc,W_{L})\in S_{L,d,r}$ then
Moreover, there is the equality
$W_{L}\begin{pmatrix}q_{\mathbf{w}_{1}}^{r}(x)\\ \vdots\\ q_{\mathbf{w}_{d}}^{r}(x)\end{pmatrix}=\begin{pmatrix}x_{1}^{r^{L-1}}\\ \vdots\\ x_{d}^{r^{L-1}}\end{pmatrix}$ (18)
where
$\begin{pmatrix}q_{\mathbf{w}_{1}}(x)\\ \vdots\\ q_{\mathbf{w}_{d}}(x)\end{pmatrix}:=(W_{L-1}\sigma_{L-2}W_{L-2}\dotsb W_{2}\sigma_{1}W_{1})(x).$
Since the polynomial neural network $(x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}})$ only has the trivial zero, Proposition 21 implies that $W_{L}$ is invertible. Multiplying both sides of (18) by $W_{L}^{-1}$, each $q_{\mathbf{w}_{i}}^{r}(x)$ is a linear combination of $x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}}$. For $r>1$, this is only possible if each $q_{\mathbf{w}_{i}}$ is a monomial. Indeed, no linear combination of $x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}}$ is a pure $r$-th power except for scalar multiples of these individual monomials. Thus each $q_{\mathbf{w}_{i}}^{r}$ is a scalar multiple of some $x_{j}^{r^{L-1}}$ and each $q_{\mathbf{w}_{i}}$ is a scalar multiple of some $x_{j}^{r^{L-2}}$. It follows that $W_{L}$ is a scaled permutation matrix—it is an invertible matrix that sends a set of monomials to a set of scalar multiples of those monomials. Thus, we have shown (17).
By induction, we get
@@
The fiber $\Psi_{\mathbf{d},r}^{-1}(\Psi_{\mathbf{d},r}(I_{d},\dotsc,I_{d}))$ then consists of the set of multi-homogeneities and has the expected dimension $(L-1)d$, which is an upper bound on the generic dimension of a fiber. It follows from the fiber-dimension theorem [39, Theorem 1.25] that
$Ld^{2}-(L-1)d$
is a lower bound for $\dim\mathcal{V}_{\mathbf{d},r}$. As $\operatorname{edim}\mathcal{V}_{\mathbf{d},r}$ is also an upper bound and agrees with the lower bound, $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r>1$ and $\operatorname{ActThr}(\mathbf{d})=1$. ∎
@@ 4 Outlook
Universal optimal bound for the activation degree threshold
Our bound in Theorem 18 on the activation degree threshold for polynomial neural networks is not universally optimal. We proved this to be the case for equi-width networks in Theorem 22. One way to improve the universal bound of activation degree thresholds is to leverage other algebro-geometric and number theoretic results. For instance, we expect that a lemma of [17] used to prove Picard theorems for holomorphic maps into Fermat hypersurfaces will allow us to improve the bound in Theorem 18.
Theorem 22 suggests that we should be looking for architectures where the activation degree threshold significantly improves upon the universal upper bound. A direct generalization of Theorem 22 in this direction is the extension to non-increasing architectures.
Pyramidal polynomial neural networks (Non-increasing architectures)
One future direction in this research area is to determine the activation degree thresholds for non-increasing architectures $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ — those architectures satisfying $d_{0}\geq d_{1}\geq\dotsb\geq d_{L}$. By generalizing the proof of Theorem 22, one might hope to prove [28, Conjecture 5.7] that for a non-increasing architecture $\mathbf{d}$ with $d_{L}>1$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r>1$. That is, prove that a non-increasing architecture $\mathbf{d}$ has activation degree threshold less than or equal to one.
@@
Another research direction is to introduce the notion of activation degree thresholds for other linear and non-linear neural networks. Permitting negative exponents in a polynomial activation map leads to the notion of rational neural networks [8]. There is experimental evidence that it is beneficial to choose rational activation maps that are very low degree. Defining activation degree thresholds for rational networks would provide a quantitative measure to make these observations rigorous. Rational neural networks are currently being investigated by Alexandros Grosdos et al. Certain linear networks may also admit the definition of an activation degree threshold as a means to study expressiveness. Tropical rational neural networks provide a lens to study networks with piecewise linear activation maps [3, 11, 12, 45]. Such networks encompass ReLU networks; a ReLU network of fixed architecture is described by a semialgebraic set inside the parameter space of tropical rational functions. In this context, tropical rational functions provide a means to define activation degree thresholds to frame and obtain results on expressiveness. Moreover, because the dimension of the neurovariety is preserved under tropicalization, activation degree thresholds may aid in translating expressiveness results for tropical neural networks to polynomial networks.
Acknowledgments
The authors would like to thank Jiayi Li, Kaie Kubjas, and Alexandros Grosdos for their comments and feedback that led to improvements of this article.